4+3[2²-3(7-9)] I already know the answer is 34, but every time I try to solve it I get a different answer. Thankyou for taking the time to help me. Please show the step by step process so that I can figure out where I went wrong.
Answer by stanbon(75887)

4+3[2²-3(7-9)]
--------
= 4 + 3[4-3(-2)]
--------
= 4 + 3[4+6]
--------
= 4 + 30
--------
= 34
==============
Cheers,
Stan H.
=============